OpenStudy (anonymous):

y=-(x-1)^4 find x intercepts

OpenStudy (anonymous):

I know how to get y intercepts.. and I guess in general i know how to find x intercepts. just rusty at breaking this out into a quadratic...

ganeshie8 (ganeshie8):

?

ganeshie8 (ganeshie8):

to find x-intercept, put y = 0

ganeshie8 (ganeshie8):

0 = -(x-1)^4 0 = x-1 1 = x x = 1

OpenStudy (anonymous):

hmm so u dont have to simplify?

ganeshie8 (ganeshie8):

already its in factored form, simplified form :)

OpenStudy (anonymous):

so what about for y = (x-4)^3 - 2

OpenStudy (anonymous):

i guess i am confused then what to do with the exponent

OpenStudy (anonymous):

\[y=(x-4)^{3}-2\]

ganeshie8 (ganeshie8):

you wanto find x-intercept or y-intercept ?

OpenStudy (anonymous):

2 = (x-4)^3 cube root each side?

OpenStudy (anonymous):

x. i know how to find y.

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so

OpenStudy (anonymous):

\[\sqrt[3]{2}=x-4\]

OpenStudy (anonymous):

\[x=4\sqrt[3]{2}\] ??

OpenStudy (anonymous):

er 4 + cube rt of 2

ganeshie8 (ganeshie8):

thank god ;)

OpenStudy (anonymous):

awesome.. thanks for your help...

ganeshie8 (ganeshie8):

looks you've mastered these ! np :)

OpenStudy (anonymous):

what did u do when the exp was 4 in the original question.

OpenStudy (anonymous):

0 = -(x-1)^4 0 = x-1 1 = x x = 1

ganeshie8 (ganeshie8):

0 = -(x-1)^4 0 = -1(x-1)^4

ganeshie8 (ganeshie8):

since -1 is NEVER equal to 0, (x-1)^4 must be equal to 0

ganeshie8 (ganeshie8):

so, 0 = (x-1)^4

OpenStudy (anonymous):

the -1 outside of the parens u mean

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

0 = (x-1)^4 take 4th root, 0 = (x-1)

OpenStudy (anonymous):

and the 4th root of 0 is 0... i see now.

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

lol so much easier than i thought.. i had x^4 + 96x^3 and other stuff lol. thanks again

ganeshie8 (ganeshie8):

those would be easy too, just we factor HCF i guess... good luck !